[Hallc_running] [New Logentry] RC Update - 02/07/22

sjdkay at jlab.org sjdkay at jlab.org
Sat Jul 2 10:55:02 EDT 2022

Logentry Text:
An update on our progress this morning, as a reminder, we won't have a meeting today and there are none planned for the long weekend.

Physics Update

- We got beam back during swing yesterday
- We finished off the Q2 = 6.0, W = 2.40 LD2+ Left (SHMS = 16.91) setting towards the end of swing
- The aluminium dummy for this setting was done too at the start of owl
- We then made progress on the LH2 data taking for the same kinematic. We're roughly halfway through this now. My current projection is that we have ~6 hours or so left on this setting, after that we proceed to the central setting (SHMS = 14.91) for the same kinematic
- The charge goal projections in the run plan are a bit off and will be updated

Accelerator Update

- We're not up to 70 or 80uA, but we have had fairly stable 50-60 uA beam through the night. The plan is to just keep this up over the long weekend
- Due to a BCM mismatch, we may have been receiving ~10% less current than we thought. This has now been corrected. As a quick example, what swing shift thought was 57uA last night was actually 52uA. See - https://logbooks.jlab.org/entry/4009749
- As a consequence of this update, the two current values on the wall display should now match up (with a small time delay)
- There have been a few accelerator issues this morning, so we're falling a bit behind schedule

Hall Update

- We've had two HMS Q1 interlocks this morning, both have cleared without issue from the CH. Something to keep an eye on though.

Plan Going Forward

- We'll keep slowly chugging through the run plan. Compared to my projection yesterday (https://logbooks.jlab.org/entry/4009519) we're roughly 5-6 hours behind my projected progress
- I'll recalibrate my crystal ball once I have more data points. Unfortunately the LH2 rates seem to be just as low as the LD2 rates for this setting.
- I've requested a BCM calibration for Tuesday, we'll see how reliable the "high" current is over the weekend and re-asses that if needed though
- Looks like the rotation to small SHMS angles (6.79) is likely to be Tuesday or even into Wednesday unless things speed up significantly

I'll drop another mini update this evening with progress. Look for another full update tomorrow morning.

Have a great long weekend!


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4009973
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220702/c89c99ac/attachment.html>

More information about the Hallc_running mailing list